OpenStudy (helpz-q):

What is the sign of the product (3)(–3)(–2)(4)? Positive, because the products (3)(–3) and (–2)(4) are negative, and the product of two negative numbers is positive Positive, because the products (3)(–3) and (–2)(4) are positive, and the product of two positive numbers is positive Negative, because the products (3)(–3) and (–2)(4) are negative, and the product of two negative numbers is negative Negative, because the products (3)(–3) and (–2)(4) are positive, and the product of two positive numbers is negative

OpenStudy (3mar):

May I help?

OpenStudy (helpz-q):

yes please

OpenStudy (3mar):

Thank you. You may know that: (-) ×(+) = (-) (+) ×(-) = (-) (+) ×(+) = (+) (-) ×(-) = (+)

OpenStudy (helpz-q):

yes

OpenStudy (3mar):

That is very good! So before you look at the choices, can you multiply these with their signs and figure out the result?

OpenStudy (helpz-q):

I am not sure? this is quite confusing

OpenStudy (3mar):

can you do two by two?

OpenStudy (helpz-q):

Huh?

OpenStudy (3mar):

(3)(–3)=?? then (–2)(4)=??

OpenStudy (helpz-q):

is it a?

OpenStudy (helpz-q):

Is it A?

OpenStudy (sshayer):

if number of negatives signs is even in multiplication then product has positive sign. if number of negative signs is odd in multiplication then product has negative sign.

OpenStudy (3mar):

\[\Huge\color{Coral}\checkmark\]

OpenStudy (3mar):